Find the centres of the spheres. which touch at the the plane 4x+3y=47 at the point (8,5,4) and which touch the Sphere x² + y ² +z² = 1

1
Expert's answer
2022-06-06T14:15:55-0400

Find the equation of normal to the plane at "(8,5,4)" :


"\frac{x-8}{4}=\frac{y-5}{3}=\frac{z-4}{0}=k,"

where "k" is some number.

Then


"x=4k+8, y=3k+5, z=4"

Center of required sphere(s) is "C(4k+8;3k+5;4)."

Radius of required sphere(s) is


"R=\sqrt{(4k+8-8)^2+(3k+5-5)^2+(4-4)^2}"


"=\sqrt{25k^2}=5|k|."

Sphere "x^2+y^2+z^2=1" has center "(0;0;0)" and radius "1."


We will have equation:


"(4k+8-0)^2+(3k+5-0)^2+(4-0)^2"

"=(5|k|+1)^2"

"16k^2+64k+64+9k^2+30k+25+16"

"=25k^2+10|k|+1"

"94k-10|k|=-104"

"k\ge0:" "84k=-104" No solution

"k<0: 94k+10k=-104"

"k=-1"


"x=4(-1)+8=4, y=3(-1)+5=2, z=4"

The center of the sphere is "C(4, 2, 4)."